The current price of a stock is $25 per share. You have $10,000 to invest. You borrow an additional $10,000 from your broker and invest $20,000 in the stock. If the maintenance margin is 30 percent, at what price will a margin call first occur?

Select one: A. $19.71. B. $17.86. C. $9.62.
Answer #1

Initial margin = The amount we have to invest/Total amount invested

Initial margin = 10,000/20,000 = 0.5 = 50%

Maintenance margin = 30%

Trigger Price = Initial Price * 1- Initial Margin 1 - Maintenance Margin

1-0.5 Trigger Price = 25 * = 1-0.3

0.5 Trigger Price = 25 * *

Trigger Price = 17.8571428571

Margin call price = $17.86 (Option B)

Can you please upvote? Thank You :-)